# Some new cases of Zilber–Pink in $Y(1)^{3}$

### v1

@@ 3.3. Isogenous triples
of degrees $N_{0}:=\deg(f_{0})$ and $N^{\prime}_{0}:=\deg(f^{\prime}_{0})$, and we impose the analogous conditions on the bases for each $H^{1}_{DR}(\mathcal{E}_{i}/C)$. We will use the notations of Section 3.2.1.
@@
Proposition 14.
[…]
@@
Let $v\in\Sigma_{K}$ and choose $\hat{v}\in\Sigma_{\hat{K}}$ lying over $v$.
If $s^{\hat{v}\textrm{-}\mathrm{an}}\in\Delta_{v}$, there exists a $\hat{v}-$adic relation $Q_{\hat{v}}$ of degree at most $4$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i\leq 3,\ 1\leq m,n\leq 2]$ generated by the elements
@@
Moreover, for $\hat{v}\in\Sigma_{\hat{K},f}$ for which the $\mathcal{E}_{i,s,k_{v}}$ are ordinary, we can choose $Q_{\hat{v}}$ independently of $v$.
@@
The first part of the proposition follows immediately from Proposition 11 (we may ignore $\mathcal{E}_{3}$ for this part).
To prove the second part, we replace $K$ with $\hat{K}$ and $v$ with $\hat{v}$, and we suppose that $v\in\Sigma_{K,f}$ is such that the $\mathcal{E}_{i,s,k_{v}}$ are ordinary.
In this case, $L:=\operatorname{End}(\mathcal{E}_{2,s,k_{v}})\otimes_{\mathbb{Z}}\mathbb{Q}$ is an imaginary quadratic field. After possibly replacing $K$ with a finite extension (which may depend on $v$ but will not be visible in our sought relation), we may choose $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ so that the action of $L$ on $H^{1}_{v}(\mathcal{E}_{2,s})$ is diagonal. This $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ determines $\{\gamma_{v,3,s_{0}},\delta_{v,3,s_{0}}\}$ and, with respect to its image in $H^{1}_{v}(\mathcal{E}_{3,s})$, the action of $\operatorname{End}(\mathcal{E}_{3,s,k_{v}})\otimes_{\mathbb{Z}}\mathbb{Q}\cong L$ is also diagonal.
If we denote by $M_{v}$ and $M^{\prime}_{v}$ the matrices representing $f$ and $f^{\prime}$ (as above), we conclude, as in the proof of Lemma 13, that the corresponding endormorphisms $\phi$ and $\phi^{\prime}$ on $H^{1}_{v}(\mathcal{E}_{2,s})$ and $H^{1}_{v}(\mathcal{E}_{3,s})$ are represented by $d_{1}N_{0}M_{v}$ and $d_{2}N^{\prime}_{0}M^{\prime}_{v}$, respectively. Hence, we conclude that $M_{v}$ and $M^{\prime}_{v}$ are diagonal.
@@
$\Lambda_{v}$ $:=(F_{2,s})^{-1}A^{T}F_{1,s}=\Omega_{v,0}M_{v}\Omega_{v,0}^{-1};$
$\Lambda^{\prime}_{v}$ $:=(F_{3,s})^{-1}(A^{\prime})^{T}F_{1,s}=\Omega_{v,0}M^{\prime}_{v}\Omega_{v,0}^{-1}.$
Since $M_{v}$ and $M^{\prime}_{v}$ are diagonal, $\Lambda_{v}$ and $\Lambda^{\prime}_{v}$ commute, so we obtain
@@
Each entry of $\Lambda_{v}$ or $\Lambda^{\prime}_{v}$ is the evaluation at $(F_{imn}^{v\textrm{-}\mathrm{an}}(x(s)^{v\textrm{-}\mathrm{an}})_{1\leq i\leq 3,\,1\leq m,n\leq 2}$ of a polynomial of degree $4$, with coefficients in $K$, which is independent of $v$. Thus, (7) gives us the desired $v$-adic relation between evaluations of $\mathcal{G}$, independent of $v$.
Again, the fact that these relations are not in $I$ can be seen by elementary computation. ∎
4. Proving Theorems 5 and 6
In this section, we prove the main height bounds, namely, Theorems 5 and 6. Before doing so, we collect some general constructions that we will use in the proofs.
@@ 4.1. Elliptic curve schemes
As per the name, for $j_{0}\in\mathbb{A}^{1}(\overline{\mathbb{Q}})\setminus\{0,1728\}$, the fiber $\mathcal{E}_{j_{0}}$ is an elliptic curve with $j$-invariant $j(\mathcal{E}_{j_{0}})=j_{0}$. The fiber at $\infty$ of the Zariski closure of $\mathcal{E}$ in $\mathbb{P}^{1}\times\mathbb{P}^{2}$ is a nodal cubic, so the fibre at $\infty$ of the connected Néron model of $\mathcal{E}$ over $S^{\prime}:=\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1728\}$ is isomorphic to $\mathbb{G}_{m}$.
Now let $n\in\mathbb{N}$ and suppose that $C\subset S_{K}^{n}$ is a geometrically irreducible algebraic curve defined over a number field $K$. For $i\in\{1,\ldots,n\}$, we obtain elliptic schemes $\mathcal{E}_{i}\to C$ by pulling back $\mathcal{E}\to S$ along the co-ordinate projections $C\to S_{K}$. We refer to the $\mathcal{E}_{i}\to C$ as the standard elliptic curve schemes on $C$.
@@ 4.2. Covering data
Consider the situation in Section 4.1 and let $C^{\prime}$ denote a smooth compactification of $C$. After possibly replacing $C^{\prime}$ with a finite étale cover, we may assume that the connected Néron model $\mathcal{E}^{\prime}_{i}$ of $\mathcal{E}_{i}$ over $C^{\prime}$ is semiabelian [Sil86, Prop. VII.5.4].
Let $s_{0}\in C^{\prime}(K)$. By [DO23, Lemma 6.6], after possibly replacing $K$ with a finite extension, we obtain
  • (a)
    a smooth projective geometrically irreducible algebraic curve $\tilde{C}^{\prime}$ over $K$;
  • (b)
    a non-constant morphism $\nu:\tilde{C}^{\prime}\to C^{\prime}$ over $K$;
  • (c)
@@
  • (d)
    a regular $\mathcal{O}_{K}$-model $\tilde{\mathfrak{C}}^{\prime}$ of $\tilde{C}^{\prime}$;
  • (e)
    semiabelian schemes $\mathfrak{G}^{\prime}_{i}\to\tilde{\mathfrak{C}}^{\prime}$
@@
  • (i)
    every point $s\in\tilde{C}^{\prime}(\overline{K})$ satisfying $x(s)=0$ is a simple zero of $x$ and satisfies $\nu(s)=s_{0}$;
  • (ii)
    $x:\tilde{C}^{\prime}\to\mathbb{P}^{1}$ is Galois, which is to say, $\operatorname{Aut}_{x}(\tilde{C}^{\prime})=\{\sigma\in\operatorname{Aut}(\tilde{C}^{\prime}):x\circ\sigma=x\}$ acts transitively on each $\overline{K}$-fibre of $x$;
  • (iii)
    $\mathfrak{G}^{\prime}_{i,K}\cong\tilde{\mathcal{E}}^{\prime}_{i}:=\tilde{C}^{\prime}\times_{C^{\prime}}\mathcal{E}^{\prime}_{i}$.
@@
By [DO23, Lemma 6.2], there exists another regular $\mathcal{O}_{K}$-model $\tilde{\mathfrak{C}}^{\prime\prime}$ of $\tilde{C}^{\prime}$ such that the $\sigma_{j}$ extend to morphisms $\tilde{\mathfrak{C}}^{\prime\prime}\to\tilde{\mathfrak{C}}$$^{\prime}$. Let $\tilde{C}=\nu^{-1}(C)$. We refer to
$\mathcal{D}:=(\mathfrak{G}^{\prime}_{1},\ldots,\mathfrak{G}^{\prime}_{n},\tilde{\mathfrak{C}}^{\prime\prime},\tilde{\mathfrak{C}}^{\prime},\tilde{C}^{\prime},\tilde{C},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
as a covering datum for $(C^{\prime},s_{0})$.
@@ 4.3. Pullback representatives
Continue from the situation obtained in Section 4.2. For $I\subset\{1,\ldots,n\}$ and $\lambda=(j,i)\in\{1,\dotsc,\ell\}\times I$, we write $\mathfrak{G}^{\prime}_{\lambda}:=\sigma^{*}_{j}\mathfrak{G}^{\prime}_{i}$ and we write $\mathcal{E}^{\prime}_{\lambda}:=\mathfrak{G}^{\prime}_{\lambda,K}$. Then $\mathfrak{G}^{\prime}_{\lambda}$ is a scheme over $\tilde{\mathfrak{C}}^{\prime\prime}$ and $\mathcal{E}^{\prime}_{\lambda}$ is a scheme over $\tilde{\mathfrak{C}}^{\prime\prime}_{K}=\tilde{C}^{\prime}$.
Let $\bar{\eta}$ denote a geometric generic point of $\tilde{C}^{\prime}$ and define an equivalence relation $\sim$ on $\{1,\dotsc,\ell\}\times I$ by the condition
@@
Let $\Lambda$ denote a set of representatives for the induced equivalence classes. We refer to $\Lambda$ as a set of pullback representatives for $(\mathcal{D},I)$. We will write $[j,i]$ for the unique element of $\Lambda$ equivalent to $(j,i)$.
@@ 4.4. Neighbourhood systems
Continue from the situation obtained in Section 4.3. To simplify notation, relabel $C^{\prime}:=\tilde{C}^{\prime}$. For each archimedean (resp. non-archimedean) place $v$ of $K$, apply [DO22, Lemma 5.5] (resp. Proposition 21) to $C^{\prime}$ and $x\in K(C^{\prime})$, thereby obtaining real numbers $r_{v}>0$ and open subspaces $U_{v,1},\dotsc,U_{v,\ell}$ of $C^{\prime{v\textrm{-}\mathrm{an}}}$ with the properties (i)–(v) of [DO22, Lemma 5.5] (resp. Proposition 21). Let $\zeta\in K^{\times}$ such that $|\zeta|_{v}\leq r_{v}$ for the finitely many places $v$ of $K$ for which $r_{v}<1$ and let $H$ denote the G-function $\zeta/(\zeta-X)$. We refer to $((r_{v},U_{v,1},\dotsc,U_{v,\ell})_{v},H)$ as a neighbourhood system for $(C^{\prime},x)$.
4.5. Proof of Theorem 5
Consider the situation in Theorem 5 and let $K$ denote the field of definition of $C$.

### v2

@@ 3.3. Isogenous triples
of degrees $N_{0}:=\deg(f_{0})$ and $N^{\prime}_{0}:=\deg(f^{\prime}_{0})$, and we impose the analogous conditions on the bases for each $H^{1}_{DR}(\mathcal{E}_{i}/C_{0})$, with constants $\mu,\mu^{\prime}\in\mathbb{Q}$, corresponding to $f_{0}$ and $f^{\prime}_{0}$, respectively. We will use the notation of Section 3.2.1.
@@
Proposition 3.6.
Let $s\in C(L)$, with $L$ a finite extension of $K$, such that there exist isogenies
@@
Let $v\in\Sigma_{K}$ and choose $w\in\Sigma_{L}$ lying over $v$.
If $s^{w\textrm{-}\mathrm{an}}\in\Delta_{v}$, then there exists a $w$-adic relation $Q_{w}$ of degree at most $4$ between the evaluations of the elements of $\mathcal{G}$ at $x(s)$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i\leq 3,\ 1\leq m,n\leq 2]$ generated by the elements
@@
Moreover, for $w\in\Sigma_{L,f}$ for which the $\mathcal{E}_{i,s,k_{v}}$ are ordinary, we can choose $Q_{w}$ independently of $v$. Otherwise, we can choose $Q_{w}$ of degree at most $2$.
@@
The existence of $Q_{w}$ with $\deg(Q_{w})\leq 2$ for $w\not\in\Sigma_{L,f}$ follows immediately from Proposition 3.2 (we may ignore $\mathcal{E}_{3}$ for these parts of the proposition).
It remains to prove that, for those $w\in\Sigma_{L,f}$ for which $\mathcal{E}_{i,s,k_{v}}$ are ordinary, we can choose $Q_{w}$ independently of $w$ and of degree at most $4$. Consider such a place $w$. Replace $K$ with $L$ and $v$ with $w$.
In this case, $E:=\operatorname{End}(\mathcal{E}_{2,s,k_{v}})\otimes_{\mathbb{Z}}\mathbb{Q}$ is an imaginary quadratic field. After possibly replacing $K$ with a finite extension (which may depend on $v$ but will not be visible in our sought relation), we may choose $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ so that the action of $E$ on $H^{1}_{v}(\mathcal{E}_{2,s})$ is diagonal. This $\{\gamma_{v,2,s_{0}},\delta_{v,2,s_{0}}\}$ determines $\{\gamma_{v,3,s_{0}},\delta_{v,3,s_{0}}\}$ via the conditions
$\gamma_{v,1,s_{0}}=f_{0}^{*}\gamma_{v,2,s_{0}}=f_{0}^{\prime*}\gamma_{v,3,s_{0}},\quad\delta_{v,1,s_{0}}=f_{0}^{*}\delta_{v,2,s_{0}}=f_{0}^{\prime*}\delta_{v,3,s_{0}}.$
The composed morphism
$H^{1}_{v}(\mathcal{E}_{2,s})\xrightarrow{\iota_{2}^{-1}}H^{1}_{v}(\mathcal{E}_{2,s_{0}})\xrightarrow{f_{0}^{*}}H^{1}_{v}(\mathcal{E}_{1,s_{0}})\xrightarrow{(f_{0}^{\prime*})^{-1}}H^{1}_{v}(\mathcal{E}_{3,s_{0}})\xrightarrow{\iota_{3}}H^{1}_{v}(\mathcal{E}_{3,s})$
conjugates the action of $E=\operatorname{End}(\mathcal{E}_{2,s,k_{v}})\otimes_{\mathbb{Z}}\mathbb{Q}$ on $H^{1}_{v}(\mathcal{E}_{2,s,k_{v}})$ into the action of $\operatorname{End}(\mathcal{E}_{3,s,k_{v}})\otimes_{\mathbb{Z}}\mathbb{Q}\cong E$ on $H^{1}_{v}(\mathcal{E}_{3,s})$. Hence the latter action is also diagonal, with respect to the basis $\{\gamma_{v,3,s_{0}},\delta_{v,3,s_{0}}\}$.
If we denote by $A$, $A^{\prime}$, $M_{v}$ and $M^{\prime}_{v}$ the matrices representing $f$ and $f^{\prime}$ (as above), we conclude, as in the proof of Lemma 3.5, that the corresponding endormorphisms $\phi$ and $\phi^{\prime}$ on $H^{1}_{v}(\mathcal{E}_{2,s})$ and $H^{1}_{v}(\mathcal{E}_{3,s})$ are represented by $d_{1}N_{0}M_{v}$ and $d_{1}N^{\prime}_{0}M^{\prime}_{v}$, respectively ($d_{1}$ as defined in Lemma 3.3). Hence, we conclude that $M_{v}$ and $M^{\prime}_{v}$ are diagonal.
@@
$\Lambda_{v}$ $:=\mu(F_{2,s})^{\rm adj}A^{T}F_{1,s}=\det(F_{2,s})\theta_{v,2}M_{v}\theta_{v,2}^{-1};$
$\Lambda^{\prime}_{v}$ $:=\mu^{\prime}(F_{3,s})^{\rm adj}(A^{\prime})^{T}F_{1,s}=\det(F_{3,s})\theta_{v,3}M^{\prime}_{v}\theta_{v,3}^{-1}.$
Since $M_{v}$ and $M^{\prime}_{v}$ are diagonal and $\theta_{v,3}=\mu(\mu^{\prime})^{-1}\theta_{v,2}$, the matrices $\Lambda_{v}$ and $\Lambda^{\prime}_{v}$ commute. In particular,
@@
Each entry of $\Lambda_{v}$ or $\Lambda^{\prime}_{v}$ is the evaluation at $(F_{imn}^{v\textrm{-}\mathrm{an}}(x(s)^{v\textrm{-}\mathrm{an}})_{1\leq i\leq 3,\,1\leq m,n\leq 2}$ of a polynomial of degree $4$, with coefficients in $K$, which is independent of $v$. Thus, (12) gives us the desired $v$-adic relation between evaluations of $\mathcal{G}$ at $s$, independent of $v$.
Again, the fact that these relations are not in $I$ can be seen by elementary computation. For example, since $A$ and $A^{\prime}$ are invertible, we can find $\gamma_{2},\gamma_{3}\in\operatorname{SL}_{2}(K)$ such that
$\gamma_{2}^{\rm adj}A^{T}=\left(\begin{matrix}\det(A)&1\\ 0&1\end{matrix}\right)\text{ and }\,\gamma_{3}^{\rm adj}(A^{\prime})^{T}=\left(\begin{matrix}\det(A^{\prime})&0\\ 1&1\end{matrix}\right).$
Define $\gamma_{1}$ to be the identity matrix so, in particular, we have
$\det(\gamma_{1})=\det(\gamma_{2})=\det(\gamma_{3}).$
Hence, $(\gamma_{1},\gamma_{2},\gamma_{3})\in V(I)$. However, we have
$\Lambda_{v}(\gamma_{1},\gamma_{2},\gamma_{3})$ $=\mu\gamma_{2}^{\rm adj}A^{T}\gamma_{1}=\mu\left(\begin{matrix}\det(A)&1\\ 0&1\end{matrix}\right),$
$\Lambda^{\prime}_{v}(\gamma_{1},\gamma_{2},\gamma_{3})$ $=\mu^{\prime}\gamma_{3}^{\rm adj}(A^{\prime})^{T}\gamma_{1}=\mu^{\prime}\left(\begin{matrix}\det(A^{\prime})&0\\ 1&1\end{matrix}\right).$
Thus, $(\gamma_{1},\gamma_{2},\gamma_{3})$ does not satisfy (12). ∎
4. Proving Theorems 1.7 and 1.8
In this section, we prove the main height bounds, namely, Theorems 1.7 and 1.8. Before doing so, we collect some general constructions that we will use in the proofs.
@@ 4.1. Elliptic curve schemes
As per the name, for $j_{0}\in S(\overline{\mathbb{Q}})$, the fiber $\mathcal{E}_{j_{0}}$ is an elliptic curve with $j$-invariant $j(\mathcal{E}_{j_{0}})=j_{0}$. The fiber at $\infty$ of the Zariski closure of $\mathcal{E}$ in $\mathbb{P}^{1}\times\mathbb{P}^{2}$ is a nodal cubic, so the fibre at $\infty$ of the connected Néron model of $\mathcal{E}$ over $\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1728\}$ is isomorphic to $\mathbb{G}_{m}$.
Now let $n\in\mathbb{N}$ and suppose that $C_{0}\subset S_{K}^{n}$ is a geometrically irreducible algebraic curve over a number field $K$. For $i\in\{1,\ldots,n\}$, we obtain elliptic schemes $\mathcal{E}_{i,0}\to C_{0}$ by pulling back $\mathcal{E}_{K}\to S_{K}$ along the co-ordinate projections $C_{0}\to S_{K}$. We refer to the $\mathcal{E}_{i,0}\to C_{0}$ as the standard elliptic curve schemes on $C_{0}$.
@@ 4.2. Covering data
Consider the situation in Section 4.1 and let $C$ denote a smooth projective model of $C_{0}$. After possibly replacing $K$ with a finite extension and $C$ with a finite étale cover, we may assume that the connected Néron model $\mathcal{E}_{i}$ of $\mathcal{E}_{i,0}$ over $C$ is semiabelian [BLR90, Sec. 7.4, Thm. 1].
Let $s_{0}\in C(K)$. By [DO23, Lemma 6.5], after possibly replacing $K$ with a finite extension, we obtain
  1. (a)
    a smooth projective geometrically irreducible algebraic curve $\tilde{C}$ over $K$;
  2. (b)
    a non-constant morphism $\nu:\tilde{C}\to C$ over $K$;
  3. (c)
    a projective regular $\mathcal{O}_{K}$-model $\mathfrak{C}$} of $\tilde{C}$;
  4. (d)
    semiabelian schemes $\mathfrak{G}_{i}\to{\mathfrak{C}}$ for $i\in\{1,\dotsc,n\}$;
  5. (e)
@@
  1. (i)
    for each $i\in\{1,\ldots,n\}$, $\mathfrak{G}_{i,K}\cong\tilde{\mathcal{E}}_{i}:=\tilde{C}\times_{C}\mathcal{E}_{i}$;
  2. (ii)
    every point $s\in\tilde{C}(\overline{K})$ satisfying $x(s)=0$ is a simple zero of $x$ and satisfies $\nu(s)=s_{0}$;
  3. (iii)
    $x:\tilde{C}\to\mathbb{P}^{1}$ is Galois, which is to say, $\operatorname{Aut}_{x}(\tilde{C})=\{\sigma\in\operatorname{Aut}(\tilde{C}):x\circ\sigma=x\}$ acts transitively on each $\overline{K}$-fibre of $x$.
@@
By [DO23, Lemma 6.2], there exists another regular projective $\mathcal{O}_{K}$-model $\mathfrak{C}$$^{\prime}$ of $\tilde{C}$ such that the $\sigma_{k}$ extend to morphisms $\mathfrak{C}$$^{\prime}\to{\mathfrak{C}}$. Let $\tilde{C}_{0}=\nu^{-1}(C_{0})$. We refer to
$\mathcal{D}:=(\mathfrak{G}_{1},\ldots,\mathfrak{G}_{n},{\mathfrak{C}}^{\prime},{\mathfrak{C}},\tilde{C},\tilde{C}_{0},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
as a covering datum for $(C,s_{0})$.
@@ 4.3. Pullback representatives
Continue from the situation obtained in Section 4.2. Fix a set of indices $I\subset\{1,\dotsc,n\}$ (this is for convenience, enabling us to have a common set-up for the proof of Theorem 1.7, which involves two indices, and the proof of Theorem 1.8, which involves three indices). For $\lambda=(k,i)\in\{1,\dotsc,\ell\}\times I$, we write $\mathfrak{G}_{\lambda}:=\sigma^{*}_{k}\mathfrak{G}_{i}$ and we write $\mathcal{E}_{\lambda}:=\mathfrak{G}_{\lambda,K}$. Then $\mathfrak{G}_{\lambda}$ is a scheme over $\mathfrak{C}$$^{\prime}$ and $\mathcal{E}_{\lambda}$ is a scheme over ${\mathfrak{C}}^{\prime}_{K}=\tilde{C}$.
Let $\bar{\eta}$ denote a geometric generic point of $\tilde{C}$ and define an equivalence relation $\sim$ on $\{1,\dotsc,\ell\}\times I$ by the condition
@@
Let $\Lambda$ denote a set of representatives for the induced equivalence classes. We refer to $\Lambda$ as a set of pullback representatives for $(\mathcal{D},I)$. We will write $[k,i]$ for the unique element of $\Lambda$ equivalent to $(k,i)$.
@@ 4.4. Neighbourhood systems
Continue from the situation obtained in Section 4.3. To simplify notation, relabel $C:=\tilde{C}$. For each archimedean (resp. non-archimedean) place $v$ of $K$, apply [DO25, Lemma 5.5] (resp. this paper’s Proposition A.3) to $C$, $\mathfrak{C}$ and $x\in K(C)$, thereby obtaining real numbers $r_{v}>0$ and simply connected (resp. simple) open subspaces $U_{v,1},\dotsc,U_{v,\ell}$ of $C^{v\textrm{-}\mathrm{an}}$ with the properties (i)–(v) of [DO25, Lemma 5.5] (resp. Proposition A.3). We refer to $\mathcal{S}:=((r_{v},U_{v,1},\dotsc,U_{v,\ell})_{v})$ as a neighbourhood system for $(C,x)$.
Let $\zeta\in K^{\times}$ such that $|\zeta|_{v}\leq r_{v}$ for the finitely many places $v$ of $K$ for which $r_{v}<1$ and let $H$ denote the G-function $\zeta/(\zeta-X)$. We refer to $H$ as an auxillary G-function for $\mathcal{S}$.
4.5. Proof of Theorem 1.7
Consider the situation in Theorem 1.7 and let $K$ denote the field of definition of $C\subset Y(1)^{3}$. To match the notation of the preceding sections, we relabel $C$ and $\overline{C}$ from the statement of Theorem 1.7 as $C_{0}$ and $C$, respectively. We break the proof into sections.